OpenStudy (anonymous):

A cell phone store sells the basic model of a phone for $30 and the upgraded model for $100 when a customer signs a two-year agreement. On Tuesday, the store has $900 in sales. The store pays $5 for each basic model and $25 for each upgraded model. If the cost of the cell phones sold on Tuesday is $200, how many upgraded models were sold? 6 phones 10 phones 15 phones 42 phones

OpenStudy (jhannybean):

$200 dollar phone? Merry Christmas.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol. ;-)

OpenStudy (dtan5457):

You need to set up 2 equations for this, a system to be exact let x represent the cheaper phone and y represent the upgraded phone for store costs you will have 5x+25y=200 for selling prices you will have 30x+100y=900 set these two equations together 5x+25y=200 30x+100y=900 multiply the first one by -4 and add them -20x-100y=-800 30x+100y=900 10x=100 x=10 10 cheaper phones, plug into any equation again 5(10)+25(y)=200 y=6<<<6 upgraded phones

OpenStudy (anonymous):

Thank you!